You have already traveled 2.3 miles. You continue to travel 0.1 miles per. minute. The expression 0.1m + 2.3 gives the distance traveled in miles after m minutes. What is miles traveled after 25 minutes and after 30 minutes.
I know the answer to 25 minutes is 4.8 but I don't know how I came to this? I am mother of child trying to solve this for 3 days. I just need an explanation of how to get the answer to this very hard question so I understand how to explain to my child. Thank you...

Answer by ikleyn(52908)

.
The expression 0.1m + 2.3 gives the distance traveled in miles after m minutes.
What is miles traveled after 25 minutes ?
~~~~~~~~~~~~~~~~~~~~~~~~~~~

Do as the condition says: substitute m= 25 into this expression 

    Distance = 0.1*m + 2.3.


You will get 0.1*25 + 2.3 = 2.5 + 2.3 = 4.8 miles.



Next, make the same for the other value m = 30:

    Distance = 0.1*30 + 2.3.


You will get 0.1*30 + 2.3 = 3 + 2.3 = 5.3 miles.

That's all.